# Uncertainty-driven Adaptive Exploration

### v2

@@ Abstract
Adaptive exploration methods propose ways to learn complex policies via alternating between exploration and exploitation. An important question for such methods is to determine the appropriate moment to switch between exploration and exploitation and vice versa. This is critical in domains that require the learning of long and complex sequences of actions. In this work, we present a generic adaptive exploration framework that employs uncertainty to address this important issue in a principled manner. Our framework includes previous adaptive exploration approaches as special cases. Moreover, we can incorporate in our framework any uncertainty-measuring mechanism of choice, for instance mechanisms used in intrinsic motivation or epistemic uncertainty-based exploration methods. We experimentally demonstrate that our framework gives rise to adaptive exploration strategies that outperform standard ones across several MuJoCo environments.
@@ 1 Introduction
Over the past decade, deep reinforcement learning (DRL) [14, 19, 32] has become the policy learning method of choice in a variety of domains. In online DRL, the goal is to approximate the $Q$-function via alternating between generating new experiences by interacting with the environment (exploration) and updating the strategy using this experience (learning) [23].
@@
Now, robotics is considered a standard domain for testing DRL [10, 13, 14, 18]. In general, such domains are characterized by: (i) continuous and multi-dimensional action space; (ii) informative, multi-component, and dense reward function (iii) episodes that are very probable to be terminated early (or “truncated”) when an agent is set in an unsafe state; (iv) transitions being highly deterministic. In such domains, the main aim is to learn a complex and long trajectory, necessitating the execution of a sequence of actions that will lead the agent to a specific state.
@@
In domains with the aforementioned characteristics, it is critical for an agent to pose the question: “Is it always better to search for new states, or should I carefully choose when to explore?” To further understand the importance of correctly answering that question, let us consider the classic Walker domain [31]. In this domain, a two-dimensional bipedal robot aims to walk in the forward direction. At the beginning of an episode, the agent is set to an initial state. An episode is truncated when the agent finds itself in an unsafe state. Suppose also that the agent has already learned an incomplete and close-to-optimal trajectory which, in this case, corresponds to performing a single step. This trajectory can guide the agent from the initial state to a new safe one. The agent’s goal is to extend this trajectory as much as possible. For that to happen, two phases are required: (i) the exploitation of the current policy by replicating the trajectory until its end (by replicating in other words the “single step” trajectory); and (ii) the exploration from the end of the trajectory until finding the next best action (explore after performing a “single step”). We should note that the switch between the two phases should happen within each episode. In case the agent explores instead of replicating the known trajectory, it is more probable to transition to a worse—and maybe unsafe—state rather than the one included in the trajectory. Given this example,  Pislar et al. [23] argue that it is more advantageous to ask “when to explore” rather than “how to explore”. [^1] The latter question is in fact the question that deep exploration methods [3, 15, 17, 21, 25] focus on. Even though these were designed for use in structurally different domains (e.g., video game domains with sparse, discrete rewards, and stochastic transitions), they are regularly used in robotics as well ([17, 18]).
@@
In this paper, we put forward ADEU - ADaptive Exploration via Uncertainty, a generic exploration framework that is able to decide in a principled manner when the switching between exploration and exploitation and vice versa should take place within an episode. Intuitively, ADEU enables exploration when it adjudges that its uncertainty regarding its policy is “high enough”. Now, uncertainty is arguably a subjective term and is highly dependent on the formulation of the problem being solved. ADEU takes that into account. It can in fact incorporate any mechanism for measuring uncertainty in its action selection mechanism. Intuitively, such an uncertainty-measuring mechanism corresponds to a particular “uncertainty definition”. Moreover, ADEU does not resort to the use of heuristic techniques or thresholds to decide the aforementioned “is uncertainty high enough” question: instead, it selects an action in a principled manner, via sampling from a distribution whose spread is proportional to its uncertainty regarding the policy. By being able to incorporate any uncertainty-measuring mechanism, and by selecting an action in the aforementioned principled manner, ADEU is in fact able to generalize over existing adaptive exploration strategies.
@@
To test our framework, we create $3$ different instances of ADEU, each one corresponding to $3$ different uncertainty-measuring mechanisms. We test our method in the most complex/difficult among the robotic tasks in [31]. Our experiments demonstrate that the performance of an ADEU instance depends on the ability of each uncertainty-measuring mechanism to quantify uncertainty. However, in our experiments, the instance of ADEU employing a specific uncertainty-measuring mechanism, outperforms the original exploration method that employs the corresponding uncertainty-measuring mechanism. Moreover, there is always some instance of ADEU that performs better than all of its opponents. As such, ADEU can be considered as a ‘plug-and-play’ adaptive exploration framework, that can be coupled with any uncertainty-measuring mechanism; and exhibits the ability to improve the performance of DRL algorithms that employ the corresponding uncertainty-measuring mechanism in their exploration process.
@@ A Motivating Example:
Let us consider a variation of the classic Frozen Lake example as shown in Figure 1.222The official documentation of the Frozen Lake environment can be found at https://gymnasium.farama.org/environments/toy_text/frozen_lake In that environment, the agent has to discover the sole trajectory that will eventually lead it to the goal passing through frozen tiles, while if it steps on a “hole” tile the episode terminates. In our example, the agent is rewarded when it reaches the next frozen tile. This reward is inversely proportional to the distance of the goal: the reward function outputs a higher reward when the agent reaches an allowed tile that is also closer to the goal. In addition, let us assume that an agent has already been trained to reach state $61$ (state at the end of the red line in Figure 1—right above state $71$). In other words, its policy $\pi(s)$ can successfully lead it to state $61$, replicating the red trajectory. The agent aims to extend the red trajectory until reaching the treasure.
Figure 1: The Frozen Lake environment, in which there is a single path connecting the initial state with the target. Exploration becomes more challenging when increasing the grid dimensions.
An effective way for that agent to explore this environment is by “trusting” its policy until the state $61$ and then switching to exploration. Assume an agent employing a classic exploration technique like $\epsilon$-greedy (or any noise-based strategy) that explores every episode with a constant $\epsilon$ (or constant variance in the case of noise-based strategy). By not modifying $\epsilon$ during an episode, the agent explores with a constant rate. Then, the agent will probably not be able to replicate the red trajectory since it is probable (approximately $\epsilon$% at each timestep) that the agent will deviate from the known policy.
In an even worse scenario, assume an exploration strategy that forces the agent to visit states with high uncertainty (e.g., a state that has not been visited enough times, or a state whose value the agent is unsure about). In that scenario, the agent is uncertain for all the states not belonging to the trajectory. Thus, this exploration strategy will force the agent to needlessly explore rather than trusting the trajectory, leading to premature termination of large number of episodes before managing to actually extend the trajectory. The use of a greedy-in-the-limit exploration strategy would lead to similar “over-exploration” phenomena.
@@
Let us define $f(s)=\frac{1}{\beta\sqrt{n(s)}}$ where $n(s)$ is a (pseudo-)counter that counts (or approximates in larger environments) the times that a state $s$ has already been visited and $\beta<1$ a positive constant. Let us also suppose that the states of the trajectory have been visited numerous times. Then, $f(s)$ arguably provides an effective signal on whether to switch (or alternate) between the two phases. This indicates that an agent should explore proportionally to the value of $f(s)$.
@@ 2.1 Formal Framework Description
Against this background, we put forward ADEU, a framework allowing for adaptive exploration, based on the uncertainty associated with each state. ADEU makes use of some uncertainty-measuring mechanism $f(s)$, whose value dictates when an agent should start exploring and when it should trust the gained knowledge, similarly to the example. As discussed earlier, uncertainty is a subjective term; its definition is closely related to the problem being solved. The ADEU framework implicitly accepts any definition of uncertainty via the implementation of an $f(s)$. As such, practically all uncertainty-measuring mechanisms—e.g., those utilized by intrinsic motivation to detect novel states, or by epistemic uncertainty-driven exploration to quantify the convergence degree of a policy—are compatible with ADEU. We will elaborate on this after formally presenting the key ADEU principles.
@@
$a(s)\sim\mathcal{D}\left(\pi(s),g(f(s))\right)$ (1)
In more detail, this action will be chosen by sampling from a distribution $D$. The first moment of that distribution is the policy $\pi(s)$ that the DRL agent has already learned. Hence, the mean of the probability distribution $D$ (intuitively, its center of mass) is the policy $\pi(s)$. The second moment (variance) of that distribution is $g(f(s))$, which is proportional to the uncertainty. Hence, $g(f(s))$ defines how “spread” the distribution is, and how close the sampled action will be to the policy $\pi(s)$.
The key intuition in ADEU is that the spread of the distribution should be defined so that it corresponds to the agent’s uncertainty regarding the $\pi(s)$ policy; while that uncertainty can be calculated in any way considered appropriate. [^3] In this paper, we use the first raw moment (mean) and the second central moment (variance), both in the context of continuous and discrete distributions; but one could conceivably use any other moments of choice to enhance the action selection. For instance, one could use some function of the third standardized moment in order to guide exploration towards a specific direction, e.g., a direction that is deemed safe [33, 34], if such a direction can be in some way associated with the kurtosis of the distribution. To this end, $f(s)$ in $g(f(s))$ is some mechanism able to compute uncertainty, like the ones used in uncertainty-based exploration methods[4, 13, 21] or mechanisms designed for intrinsic reward construction in intrinsic motivation methods[1, 3, 17]. [^4] To clarify, in our method, we use the mechanism used in intrinsic motivation methods for defining $f(s)$, and not for providing an exploration bonus. By default, $f(s)$ will be outputting a positive scalar. In multidimensional distributions $D$, this scalar can multiply an Identity matrix. In addition, since $f(s)$ can be any uncertainty-measuring mechanism, we use $g(\cdot)$ as a normalizer to prevent $f(s)$ from outputting enormous values.
@@
Note that the discussion above does not imply that ADEU needs to be provided with any trajectories or other forms of background knowledge. ADEU assumes a zero-length initial trajectory at the beginning of the training process, and then is progressively able to extend it via its adaptive exploration technique. [^5] Of course, if ADEU is provided with an accurate initial trajectory or policy, it can exploit such knowledge. In other words, at the start of training, the agent has not yet converged at a policy, even for the initial states. As such, it lacks certainty about which actions to take and must explore until it converges to a reliable policy even for the initial state. Once this occurs, the agent follows the learnt action policy from the initial state until it reaches the next state where “enough” uncertainty arises. There, it resumes exploration. This cycle of exploration and policy refinement continues until the agent becomes confident across the entire trajectory.
@@
ADEU’s efficiency arises from the fact that it does not needlessly explore, but instead replicates the parts of a “known” or “learnt” trajectory that it trusts. But this trajectory might be sub-optimal. So, to keep improving the learnt trajectory, ADEU also performs the so-called rollout episodes. In more detail, at the beginning of every episode, with probability $\rho$ ADEU performs exploration using as uncertainty mechanism the $f_{\text{rollout}}(s)=c$, where $c$ is a constant number. Hence, for that specific “rollout” episode, ADEU takes the chance to explore so as to improve the learnt trajectory. We should note that in the case of a well-constructed $f(s)$ that smoothly decreases, $\rho$ should and can simply be set to 0. For simplicity, in our implementation we treat $\rho$ as a constant that does not change throughout the training process. Also, we set $f_{\text{rollout}}(s)=c$ to present a simple version of our framework, but, in more complex environments, this can change too.

### v3

@@ Abstract
Adaptive exploration methods learn complex policies via alternating between exploration and exploitation. An important question for such methods is to determine the appropriate moment to switch between exploration and exploitation and vice versa. This is critical in domains that require the learning of long and complex sequences of actions. In this work, we present a generic adaptive exploration framework that employs uncertainty to address this important issue in a principled manner. Our framework includes previous adaptive exploration approaches as special cases. Moreover, it can incorporate any uncertainty-measuring mechanism of choice, such as mechanisms used in intrinsic motivation, or epistemic uncertainty-based exploration methods; and is experimentally shown to give rise to adaptive exploration strategies that outperform standard ones across several environments. Moreover, we showcase its potential for utilization in safety-critical domains.
@@ 1 Introduction
Over the past decade, deep reinforcement learning (DRL) [Lillicrap2016ddpg, Mnih2015dqn, van2016ddqn] has become the policy learning method of choice in a variety of domains. In online DRL, the goal is to approximate the $Q$-function via alternating between generating new experiences by interacting with the environment (exploration) and updating the strategy using this experience (learning) [pislar2022when].
@@
Now, robotics is considered a standard domain for testing DRL [fortunato2018noisy, lee2021sunrise, Lillicrap2016ddpg, mahankali2024random]. In general, such domains are characterized by: (i) continuous and multi-dimensional action space; (ii) informative, multi-component, and dense reward function (iii) episodes that are very probable to early termination (or “truncation”) when an agent is set in an unsafe state; (iv) transitions being highly deterministic. In such domains, the main aim is to learn a complex and long trajectory, necessitating the execution of a sequence of actions that will lead the agent to a specific state.
@@
In domains with the aforementioned characteristics, it is critical for an agent to pose the question: “Is it always better to search for new states, or should I carefully choose when to explore?” To further understand the importance of correctly answering that question, let us consider the classic Walker domain [towers2024gymnasium]. In this domain, a two-dimensional bipedal robot aims to walk in the forward direction. At the beginning of an episode, the agent is set to an initial state. An episode is truncated when the agent finds itself in an unsafe state. Suppose also that the agent has already learned an incomplete and close-to-optimal trajectory which, in this case, corresponds to performing a single step. This trajectory can guide the agent from the initial state to a new safe one. The agent’s goal is to extend this trajectory as much as possible. For that to happen, two phases are required: (i) the exploitation of the current policy by replicating the trajectory until its end (by replicating in other words the “single step” trajectory); and (ii) the exploration from the end of the trajectory until finding the next best action (explore after performing a “single step”). We should note that the switch between the two phases should happen within each episode. In case the agent explores instead of replicating the known trajectory, it is more probable to transition to a worse—and maybe unsafe—state rather than the one included in the trajectory. Given this example, pislar2022when argue that it is more advantageous to ask “when to explore” rather than “how to explore”. [^1] The latter is in fact the question that deep exploration methods [burda2018rnd, lin2024curse_of_diversity, Lobel2023FlippingCoins, osband2016bdqn, sheikh2022maximizing_ensemble_diversity] focus on. Even though these were designed for use in structurally different domains (e.g., video game domains with sparse, discrete rewards, and stochastic transitions), they are regularly used in robotics as well [Lobel2023FlippingCoins, mahankali2024random]
@@
In this paper, we present ADEU - ADaptive Exploration via Uncertainty, a generic exploration framework that decides in a principled manner when the switching between exploration and exploitation and vice versa should take place within an episode. Intuitively, ADEU enables exploration when it adjudges that its uncertainty regarding its policy is “high enough”. Now, uncertainty is arguably a subjective term and is highly dependent on the formulation of the problem being solved. ADEU takes that into account. It can, in fact, incorporate any mechanism for measuring uncertainty in its action selection mechanism. Intuitively, such an uncertainty-measuring mechanism corresponds to a particular “uncertainty definition”. Moreover, ADEU does not resort to the use of heuristic techniques or thresholds to answer the aforementioned “is uncertainty high enough” question: instead, it selects an action in a principled manner, by sampling from a distribution whose spread is proportional to its uncertainty regarding the policy. By being able to incorporate any uncertainty-measuring mechanism, and by selecting an action in the aforementioned principled manner, ADEU is, in fact, able to generalize over existing adaptive exploration strategies. Furthermore, we prove in the paper that ADEU can achieve optimal exploratory behaviour for a wide class of games. Finally, we put forward ADEU versions that can safely explore in safety-critical environments, as confirmed by preliminary results.
@@
To test our framework, we create different instances of ADEU, each corresponding to a different uncertainty-measuring mechanism. We test our method in the most complex/difficult among the robotic tasks in [towers2024gymnasium] and in the DeepSea domain—a domain which tests the agent’s ability for deep exploration. Our experiments demonstrate that the performance of an ADEU instance depends on the ability of each uncertainty-measuring mechanism to quantify uncertainty. However, in our experiments, the instance of ADEU employing a specific uncertainty-measuring mechanism, outperforms the original exploration method that employs the same uncertainty-measuring mechanism. Moreover, there is always some instance of ADEU that performs better than all of its opponents. As such, ADEU can be considered as a ‘plug-and-play’ adaptive exploration framework, that can be coupled with any uncertainty-measuring mechanism; and exhibits the ability to improve the performance of DRL algorithms that utilize the corresponding uncertainty-measuring mechanism in their exploration process.
@@ A Motivating Example:
Let us consider a variation of the classic Frozen Lake example as shown in Figure 1.222Frozen Lake’s official documentation: https://gymnasium.farama.org/environments/toy_text/frozen_lake In that environment, the agent has to discover the sole trajectory that will eventually lead it to the goal, passing through frozen tiles, while if it steps on a “hole” tile the episode terminates. In our example, the agent is rewarded when it reaches the next frozen tile. This reward is inversely proportional to the distance of the goal: the reward function outputs a higher reward when the agent reaches an allowed tile that is also closer to the goal. In addition, let us assume that an agent has already been trained to reach state $61$ (state at the end of the red line in Figure 1—right above state $71$). In other words, its policy $\pi(s)$ can successfully lead it to state $61$, replicating the red trajectory. The agent aims to extend the red trajectory until reaching the treasure.
Figure 1: The modified Frozen Lake environment. Exploration gets harder as dimensions increase.
An effective way for that agent to explore this environment is by “trusting” its policy until state $61$ and then switching to exploration. Assume an agent employing a classic exploration technique like $\epsilon$-greedy (or any noise-based strategy) that explores every episode with a constant $\epsilon$ (or constant variance in the case of a noise-based strategy). By not modifying $\epsilon$ during an episode, the agent explores at a constant rate. Then, the agent will probably not be able to replicate the red trajectory since it is probable (approximately $\epsilon$% at each timestep) that the agent will deviate from the known policy.
In an even worse scenario, assume an exploration strategy that forces the agent to visit states with high uncertainty (e.g., a state that has not been visited enough times, or a state whose value the agent is unsure about). In that scenario, the agent is uncertain for all the states not belonging to the trajectory. Thus, this exploration strategy will force the agent to needlessly explore rather than trusting the trajectory, leading to the truncation of a large number of episodes before managing to actually extend the trajectory. The use of a greedy-in-the-limit exploration strategy would lead to similar “over-exploration” phenomena.
@@
Let us define $f(s)=\frac{1}{\beta\sqrt{n(s)}}$ where $n(s)$ is a (pseudo-)counter that counts (or approximates in larger environments) the times that a state $s$ has already been visited, and $\beta<1$ a positive constant. Let us also suppose that the states of the trajectory have been visited numerous times. Then, $f(s)$ arguably provides an effective signal on whether to switch/alternate between the two phases. This suggests an agent should explore proportionally to the value of $f(s)$.
@@ 2.1 Formal Framework Description
Against this background, we put forward ADEU, a framework allowing for adaptive exploration, based on the uncertainty associated with each state. ADEU makes use of some uncertainty-measuring mechanism $f(s)$, whose value dictates when an agent should start exploring and when it should trust the gained knowledge, similarly to the example. As discussed earlier, uncertainty is a subjective term; its definition is closely related to the problem being solved. The ADEU framework implicitly accepts any definition of uncertainty via the implementation of an $f(s)$. As such, practically all uncertainty-measuring mechanisms—e.g., those utilized by intrinsic motivation methods to detect novel states, or by epistemic uncertainty-driven exploration to quantify the convergence degree of a policy—are compatible with ADEU. We will elaborate on this after formally presenting the key ADEU principles.
@@
$a\sim\mathcal{D}\left(\pi(s),g(f(s))\right)$ (1)
where $\mathcal{D}$ is the distribution of choice, $\pi(s)$ is the current policy for a state $s$, $f(s)$ is the aforementioned uncertainty-measuring mechanism and $g()$ is a normalizer function.
In more detail, this action will be chosen by sampling from a distribution $D$. The first moment of that distribution is the policy $\pi(s)$ that the DRL agent has already learned. Hence, the mean of the probability distribution $D$ (intuitively, its center of mass) is the policy $\pi(s)$. The second moment (variance) of that distribution is $g(f(s))$, which is proportional to the uncertainty. Hence, $g(f(s))$ defines how “spread” the distribution is, and how close the sampled action $a$ will be to the policy $\pi(s)$.
The key intuition in ADEU is that the spread of the distribution should be defined so that it corresponds to the agent’s uncertainty regarding the $\pi(s)$ policy; while that uncertainty can be calculated in any way considered appropriate. [^3] In this paper, we use the first raw moment (mean) and the second central moment (variance), both in the context of continuous and discrete distributions. One could conceivably use any other moments of choice to enhance action selection. For instance, one could use some function of the third standardized moment in order to guide exploration towards a specific direction, e.g., a direction that is deemed safe [wachi2023safe, ijcai2022p3o], if such a direction can be in some way associated with the kurtosis of the distribution. To this end, $f(s)$ in $g(f(s))$ is some mechanism able to compute uncertainty, like the ones used in uncertainty-based exploration methods[Chen2017ucb, lee2021sunrise, osband2016bdqn] or mechanisms designed for intrinsic reward construction in intrinsic motivation methods[Badia2020NeverGiveUp, burda2018rnd, Lobel2023FlippingCoins]. [^4] To clarify, in our method, we use the mechanism used in intrinsic motivation methods for defining $f(s)$, and not for providing an exploration bonus. By default, $f(s)$ will output a positive scalar. In multidimensional distributions $D$, this scalar can multiply an Identity matrix. In addition, since $f(s)$ can be any uncertainty-measuring mechanism, we use $g(\cdot)$ as a normalizer to prevent $f(s)$ from outputting enormous values. Equipped with $f$ and $g$, ADEU can effectively utilize any uncertainty-measuring mechanism proposed in previous (deep exploration) works.
@@
Note that the discussion above does not imply that ADEU needs any trajectories or other forms of background knowledge. ADEU assumes a zero-length initial trajectory at the beginning of the training process, and then is progressively able to extend it via its adaptive exploration technique. [^5] Of course, if ADEU is provided with an accurate initial trajectory or policy, it can exploit such knowledge. In other words, at the start of training, the agent has not yet converged at a policy, even for the initial states. As such, it lacks certainty about which actions to take and must explore until it converges to a reliable policy, even for the initial state. Once this occurs, the agent follows the learnt action policy from the initial state until it reaches the next state where “enough” uncertainty arises. There, it resumes exploration. This exploration and policy refinement cycle continues until the agent becomes confident across the entire trajectory.
We now proceed to define a generic class of exploration-related decision problems (or exploration-related single-agent “games”) such as the “modified Frozen Lake” one, for which ADEU agents are able to achieve optimal exploratory behaviour. These are“increasing-reward” settings: intuitively, only the best action transfers the agent from state $s$ to a “better” state $s^{*}$ where $\max_{a}r(s^{*},a)\geq\max_{a}r(s,a)$. (In our definition and discussion below, $s$ is the current state, $s^{\prime}$ some future ones, and $s^{*}$ is the best future one. We will also say that $s$ “neighbours” $s^{\prime}$ if $T(s,\cdot,s^{\prime})>0$).
Definition 2.1 (Increasing-Reward Single-Agent Games).
In Increasing-Reward Single-Agent Games, an agent has to decide on an action $a$ to execute at a state $s$, given that the following hold.
  • $\exists!\langle s^{*},a^{*}\rangle:r(s^{*},a^{*})\geq\max_{a}r(s,a)$. There is a unique future $s^{*},a^{*}$ pair, with $s^{*}$ neighbouring $s$, in which the agent achieves a higher reward than when playing $\arg\max_{a}r(s,a)$ in $s$.
  • $\forall s^{\prime}\neq s^{*},\quad\exists a:max_{a^{\prime}}r(s^{\prime},a^{\prime})<r(s,a)$. That is, for all future states $s^{\prime}\neq s^{*}$ “neighbouring” $s$, the agent achieves a lower maximum reward than the reward achieved in the current state $s$ for at least one action $a$.
  • $\exists!b:b=\arg\max_{a}r(s,a),\text{ }\forall s$. That is, at every state $s$, there is a unique action $b$ that maximizes the reward function in that state $s$. Moreover, for that same $b$ it holds:
  • $T(s,b,s^{*})=1$, where $r(s^{*},a^{*})>\max_{a^{\prime}}r(s^{\prime},a^{\prime})$ where $s^{\prime}$ neighbours $s$, $s^{\prime}\neq s^{*}$, and $\langle s^{*},a^{*}\rangle$ is the unique pair mentioned above—i.e., when the agent executes $b$ in $s$, it will transition with probability $1$ to the (unique) $s^{*}$ that is (strictly) better (in terms of maximum reward) than any other neighbour state $s^{\prime}$.
  • Increasing-Reward Single Agent Games inherit all other characteristics from standard MDP models [Sutton1998RL].
Corollary 2.1.
In Increasing-Reward Single-Agent Games, the optimal behaviour for an agent is to explore a state $s$ until action $b:b=\arg\max_{a}r(s,a)$ is found; no further exploration is then required for $s$.
Notice that a plethora of standard games (such as the one described in Fig. 1, and also sorting, finding the sole existing trajectory in complex maze environments, etc.) are considered as Increasing-Reward Single-Agent Games. We now prove that ADEU (equipped with an effective $f(s)$) achieves efficient exploratory behaviour in such games with a large but finite action space. In more detail, assuming that after a finite amount of time, the ADEU agent will discover action $b$ for a state $s$, the $f(s)$ will decrease and stop the exploration in that specific state $s$. Hence, for this ADEU agent $\pi(s)=b$ and the agent will explore the future state $s^{*}$. A suitable $f(s)$ should measure uncertainty as, e.g., the difference between the current policy $\pi(s)$ and $b=\arg\max_{a}r(s,a)$; or as the $f(s)$ used in the modified Frozen Lake (the latter is described in the Appendix).
Definition 2.2.
Following the aforementioned notation, we define as “best neighbour state at timestep i1i-1” the state $s_{i}^{*}:T(s_{i-1},b,s_{i}^{*})=1$ and $b=\arg\max_{a}r(s_{i-1},a)$
Theorem 1.
ADEU agents that have executed $b:=\arg\max_{a}r(s_{i},a)$ for some $s_{i}$ during a training episode, will reduce their exploration of $s_{i}$ relative to the effectiveness of their $f(s)$ mechanism and focus on exploring $s_{i+1}^{*}$.
Proof.
To prove Theorem 1, we will use mathematical induction.
  • Base Case ($i=0$): The agent will start executing action $b_{0}$, after (randomly) exploring $s_{0}$ for a finite amount of time. In addition, it will reduce its exploration of new actions at $s(0)$ since $f(s_{0})$ decreases, thus learning $\pi(s_{0})=b$. (The frequency by which the execution of $b$ will increase and the exploration of other actions at $s$ will decrease, is relative to the effectiveness of $f(s)$). Moreover, the agent will start exploring actions at $s^{*}_{1}$ (due to transitioning there via $b$).
  • Inductive Hypothesis ($i=k$): We assume that $f(s_{j})\rightarrow 0\forall j=0...k-1$, and learned $\pi(s_{j}^{*})=b_{j}$; and that the agent has found the action $b_{k}$ at $s_{k}$. As such, it decreases exploration in state $s_{k}$. The agent will reach $s_{k+1}^{*}$ for the first time (due to transitioning there via $b_{k}$).
  • Inductive Step ($i=k+1$): Due to the inductive hypothesis, the agent will reach $s_{k+1}^{*}$ for the first time. Then, since $f(s_{k+1}^{*})>>0$ (there is high uncertainty in $s_{k+1}^{*}$, the agent will explore state $s_{k+1}^{*}$ until eventually finding $b_{k+1}$ and eventually learning $\pi(s_{k+1}^{*})=b_{k+1}$. In addition, $f(s_{k+1}^{*})$ will drop until reaching $f(s_{k+1}^{*})\rightarrow 0$, which will prompt the agent to reduce exploration in $s_{k+1}^{*}$ and start exploring $s_{k+2}^{*}$.
Theorem 1 shows that an ADEU agent, when provided with an effective uncertainty-measuring mechanism, will solve an Increasing-Reward Single-Agent Game. The theorem can be extended to guarantee efficient (in expectation) ADEU behavior in enviornments with stochastic reward or transition functions also.
@@
ADEU’s efficiency arises from the fact that it does not needlessly explore, but instead replicates the parts of a “known” or “learnt” trajectory that it trusts. But this trajectory might be sub-optimal. To keep improving the learnt trajectory, ADEU also performs the so-called rollout episodes. In more detail, at the beginning of every episode, with probability $\rho$, ADEU performs exploration using as uncertainty mechanism the $f_{\text{rollout}}(s)=c$, where $c$ is a constant number. Hence, for that specific “rollout” episode, ADEU takes the chance to explore so as to improve the learnt trajectory. We should note that in the case of a well-constructed $f(s)$ that smoothly decreases, $\rho$ should and can simply be set to 0. For simplicity, in our implementation, we treat $\rho$ as a constant that does not change throughout the training process. Also, we set $f_{\text{rollout}}(s)=c$ to present a simple version of our framework, but, in more complex environments, this can change too.
Both the rollout episodes and the potential use of more complex definitions of uncertainty, allow the ADEU agent to effectively explore different, and more complex settings (e.g., even settings requiring deep exploration like DeepSea [osband2019deepSea]) than the ones in Definition 2.1, as discussed immediately below and as our experiments in Section 3 demonstrate.